Indicate the concentration of each ion present in the solution formed by mixing the following. (Assume that the volumes are additive.) (a) 10 mL of 0.100 M HCl and 10.0 mL of 0.510 M HCl

H+

Cl -
Answer #1

solution a

millimoles of HCl = 10 x 0.100 = 1.0

solution b

millimoles of HCl =10 x 0.510 = 5.1

total millimoles = 5.1 + 1.0 = 6.1

total volume = 10 + 1 = 20.0 mL

[HCl] = 6.1 / 20 = 0.305 M

as HCl dissociates completely

[H+] = 0.305 M

[Cl-] = 0.305 M